40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 22 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  1551   Sat Jun 20 22:59:58 2015 EvanDailyProgressopticSouth Faraday isolator path reworked

I reworked the beginning of the south optical path so that there are two steering mirrors before the beam goes into the FI.

Recall that previously we had no steering mirrors before the FI. Then in December, I just moved the FI sightly downstream, so that there was one mirror before the FI.

Today I added two steering mirrors  (Y1-1025-45P) in such a way that the total path length should be more or less unchanged. The first lens after the laser is now placed after the first steering mirror. (I tried to place it so that it has the same displacement from the laser head as it did previously.) The FI is placed after the second steering mirror, and it is immediately followed by a HWP.

Ideally we would maybe put down another HWP before the FI, since the steering mirrors are only HR for p-pol, and the beam on the first two steering mirrors is some combination of s-pol and p-pol (since we use a HWP + the FI to control the power after the FI).

After steering through the FI, the beam looks pretty round on the IR card. I don't see any spray or stray beams.

I tuned the pre-FI HWP so that there is now 20.4 mW transmitted through the FI. The power transmitted through the 21.6 MHz EOM (which is after the third steering mirror) is 19.6 mW. I also don't see any spray on transmission.

Attachment 1: southfi.jpg
southfi.jpg
  1550   Sat Jun 20 10:14:50 2015 EvanNotesopticcoating optimization for AlGaAs:electric field in coating layer

I reran multidiel_rt with the as-built coating structure. The penetration depth is x0 = 560 nm. With A = 5.6 ppm absorption on each mirror, the absorption coefficient is therefore α = 0.05 cm−1.

Penetration depth x0 is defined via E(x)/E(0) = exp[−x/(2x0)]. Absorption coefficient is defined as α = A/(2x0), since the effective distance traveled through the coating is 2x0. [I belive this is the same definition that Garrett uses.]

The script for this is in the paper directory of the svn, under source files.

Attachment 1: Efieldtrans.pdf
Efieldtrans.pdf
  1549   Mon May 18 16:49:59 2015 AidanSummaryPMCRealigning PMC with rubbish input beam

I'm practicing the procedure of aligning the PMC. The input beam isn't great (I might be clipping a little on the Faraday). I'm mostly trying to get the hang of aligning into the cavity again as it's been a while.

My technique, so far, is to adjust the alignment, in yaw, of the third mirror before the cavity and then sweep the final mirror before the cavity slowly in yaw. I'm looking for flashes of transmission on the trans-camera. The goal is to keep tweaking until the order of the transitted mode is reduced to a few or zero. I should probably be sweeping the PZT at the same time.

I'm going to adjust the alignment through the FI tomorrow - just as soon as I get my hands on a couple more mirrors.

  1548   Thu Apr 16 13:31:32 2015 AidanSummaryDAQAcromag XT1541 DAC working correctly - a little trick to configure

https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:resources:computing:acromag

 

Quote:

Aidan.

Success!

I configured the Acromag XT1541 DAC to run with EPICS. This was a touch trickier than the ADC as there is a subtlety with the channel configuration in the EPICS database. The bottom line is that now I can change the value in an EPICS channel and a multimeter attached to the unit will show a corresponding change in voltage.

The attached files (ioctest2.cmd and IOCTEST2.db) are used to access the first output channel, OUT00, on the unit. Now that I've got the thing working I can debug the calibration. Once that's sorted I'll summarize the set-up procedure on a Wiki page with glorious detail for future reference.

The following command line is used to open the modbus EPICS server.

${EPICS_MODULES}/modbus/bin/${EPICS_HOST_ARCH}/modbusApp ioctest2.cmd

The ioctest1 files are for the ADC unit.

 

  1547   Wed Apr 15 17:51:43 2015 AidanSummaryDAQAcromag XT1541 DAC working correctly - a little trick to configure

Aidan.

Success!

I configured the Acromag XT1541 DAC to run with EPICS. This was a touch trickier than the ADC as there is a subtlety with the channel configuration in the EPICS database. The bottom line is that now I can change the value in an EPICS channel and a multimeter attached to the unit will show a corresponding change in voltage.

The attached files (ioctest2.cmd and IOCTEST2.db) are used to access the first output channel, OUT00, on the unit. Now that I've got the thing working I can debug the calibration. Once that's sorted I'll summarize the set-up procedure on a Wiki page with glorious detail for future reference.

The following command line is used to open the modbus EPICS server.

${EPICS_MODULES}/modbus/bin/${EPICS_HOST_ARCH}/modbusApp ioctest2.cmd

The ioctest1 files are for the ADC unit.

Attachment 1: ioctest2.cmd
epicsEnvSet("IOC", "ioctest2")
epicsEnvSet("ARCH","linux-x86_64")
epicsEnvSet("TOP","${EPICS_MODULES}/modbus")
epicsEnvSet("MDBTOP","${EPICS_MODULES}/modbus")
dbLoadDatabase("$(MDBTOP)/dbd/modbus.dbd")
modbus_registerRecordDeviceDriver(pdbbase)
drvAsynIPPortConfigure("c3test2","10.0.1.41:502",0,0,1)
modbusInterposeConfig("c3test2",0,5000,0)
drvModbusAsynConfigure("DAC_Reg","c3test2",0,6,1,8,4,0,"Acromag")
dbLoadDatabase("./IOCTEST2.db")
... 2 more lines ...
Attachment 2: IOCTEST2.db
record(ao, "C3:ACROMAG_OUTPUT0")
{
	field(DTYP, "asynInt32")
	field(OUT, "@asynMask(DAC_Reg, 0, -16)MODBUS_DATA")
   	field(LINR,"NO CONVERSION")
	field(EGUL, "-163.84")
	field(EGUF, "163.835")
	field(HOPR, "163.835")
	field(LOPR, "-163.84")
	field(PREC, "2")
... 3 more lines ...
Attachment 3: ioctest1.cmd
epicsEnvSet("IOC", "ioctest1")
epicsEnvSet("ARCH","linux-x86_64")
epicsEnvSet("TOP","${EPICS_MODULES}/modbus")
epicsEnvSet("MDBTOP","${EPICS_MODULES}/modbus")
dbLoadDatabase("$(MDBTOP)/dbd/modbus.dbd")
modbus_registerRecordDeviceDriver(pdbbase)
drvAsynIPPortConfigure("c3test1","10.0.1.42:502",0,0,1)
modbusInterposeConfig("c3test1",0,5000,0)
drvModbusAsynConfigure("ADC_Reg","c3test1",0,4,0,8,4,32,"Acromag")
dbLoadDatabase("./IOCTEST1.db")
... 2 more lines ...
Attachment 4: IOCTEST1.db
record(ai, "C3:ACROMAG_INPUT0")
{
	field(SCAN, ".1 second")
	field(FLNK, "C3:ACROMAG_INPUT0_C.PROC")
	field(DTYP, "asynInt32")
	field(INP, "@asynMask(ADC_Reg 0 -16)MODBUS_DATA")
}
record(ai, "C3:ACROMAG_INPUT1")
{
	field(SCAN, ".1 second")
... 5 more lines ...
  1546   Fri Apr 10 09:33:43 2015 EvanSummaryComputersAcromag ADC set up
Quote:

I set up an Acromag slow controls based on the procedure that Keith wrote in T1400200. It's really pretty easy. It took an hour and 15 minutes from installing Ubuntu on a machine to having a functioning ADC channel from the Acromag unit. I haven't yet set up a DAC unit - this will require some tweaking of some of the EPICS parameters. Once I've done that I'll upload a complete procedure to the Wiki.

This is relatively promising for supporting/replacing VME slow channels.

yesyes

  1545   Thu Apr 9 17:03:35 2015 AidanSummaryComputersAcromag ADC set up

I set up an Acromag slow controls based on the procedure that Keith wrote in T1400200. It's really pretty easy. It took an hour and 15 minutes from installing Ubuntu on a machine to having a functioning ADC channel from the Acromag unit. I haven't yet set up a DAC unit - this will require some tweaking of some of the EPICS parameters. Once I've done that I'll upload a complete procedure to the Wiki.

This is relatively promising for supporting/replacing VME slow channels.

  1544   Sat Apr 4 18:15:39 2015 ranaDailyProgressPMCBeware the Bow-Tie

Recall the horrible sadaftermath of using the bow-tie configurationno for the green laser ALS setup at the 40m:angry

All of the supposedly HR mirrors leaked through enough lightcrying to make the table into a Christmas tree. The bow-tie should only be allowed when using the super HR mirrors from G&H which have T < 100 ppm over a wide range of angles.cheeky

  1543   Wed Apr 1 17:33:08 2015 AidanDailyProgressPMCLaser out of alignment to PMC

I've been trying to lock the laser to the PMC since we adjusted the Faraday. It's basically badly out of alignment now. I can only see very higher order modes flash when I scan the cavity.

The problem, currently, is that the Faraday is too high and we don't have enough mirrors to control the beam going through it. I'm going to install a second mirror in a bow-tie configuration tomorrow and realign the beam through the Faraday, 21.5MHz EOM and, hopefully, the PMC.

I also spent a lot of today tracing out the control loop for the laser slow control. Once I have all the control loops understood, I'm going to draw a diagram for the Wiki.

P.S. Found out that the PSL PMC Servo board is D980352. I've updated the Electronics page on the Wiki to indicate this...

  1542   Tue Mar 31 11:43:37 2015 AidanSummaryComputersRebooted VME and added sitemap

Aidan

Aidan rebooted the Sun machine and VME. It took a while to get the EPICS channels to work again. The following seemed to work:

  1. Reboot Sun machine.
  2. Reboot the VME crate by depressing the reboot button on the top of the crate.
  3. Log into VME (psl1) at 10.0.0.2 from the Sun machine.
  4. Check the existence of various channels with dbpr "C3:PSL-RCAV_RCPID_SETPOINT"
  5. On the Sun, cd to /usr1/epics/psl/scripts and run "perl rcav_PID_2012_06_15.pl"
  6. Confirmed that PID values started updating on the Sun screen.

Aidan also added a sitemap (~/sitemap.adl), see attached image, for the CTN lab. Aidan added an alias to /home/controls/.bashrc

Aidan

The alias is:

alias sitemap="medm -x /home/controls/sitemap.adl"

Aidan

 

Attachment 1: Screenshot-sitemap.adl.png
Screenshot-sitemap.adl.png
  1541   Mon Mar 30 14:29:25 2015 AidanSummaryEnvironmentCleaning underway

I've started removing a lot of the miscellaneous hardware from the lab (old pieces of Bosch framing, sheets of acrylic/plastic, etc). Some has gone into the ATF - we'll have to decide whether to keep it permanently or not. Right now, like Indiana Jones and the Last Crusade, I'm trying to see if there truly exists some space in this lab - or if it is a myth propagated though the ages. 

  1540   Mon Mar 30 14:24:48 2015 AidanNotesComputersLab workstation has died

The Windows XP workstation in the lab died last week. After booting up to the "Windows XP" screen it reported a hardware problem. The subsequent reboot got to the BiOS and reported a corrupted memory problem. I'm going to pull the hard drive and replace the computer.

  1539   Mon Mar 30 14:22:36 2015 AidanNotesComputersOrdered Acromag slow controls test unit

I ordered the following Acromag units for a new slow controls test setup. The idea is to replace the Sun workstation running the VME slow controls.

Keith Thorne has already used some code to interface to these controls at the site:

As for Modbus, that package relies on the ASYN package, also for EPICS"
If you want it all put together for you, you can find some pre-built stuff at
 
ready to install at /ligo/apps/ubuntu12
 
source code ready to be built
 
- This has all the LIGO patches for macros, long variable names, etc. detail in 
  1538   Wed Feb 4 17:29:29 2015 AidanSummaryDocumentationElectronics and power chassis connection summary

I spent this afternoon tracing out all the connections to and from all the chassis in the CTN lab.

We currently have 19 power supplies in use. Take a moment and think about that.

Attachment 1: labbook_1.pdf
labbook_1.pdf
Attachment 2: labbook_2.pdf
labbook_2.pdf
Attachment 3: labbook_3.pdf
labbook_3.pdf
Attachment 4: labbook_1a.pdf
labbook_1a.pdf
  1537   Wed Dec 17 20:30:36 2014 Aidan, EvanDailyProgress Minor changes to layout of cavity transmitted path

Quote:

For one, we removed the QWPs which were the first optics in the transmission paths. These had been necessary for the prior cavities where the Silica Tantala mirror coatings were not birefringent. The circular polarization which was transmitted needed to be turned into linear polarization to get the beat note on the PD. Now, because the cavities with AlGaAs coatings are birefringent, the resonant and transmitted light is already linearly polarized and the QWPs unnecessary. Before removing them, the power on the main readout PD, a PD1811, was 208 mV. Afterwards, it was 194 mV.

On the south path, we have placed a HWP so that the transmitted beams can have their polarizations matched. It is on a 1" post and held down with a fork.

In the longer term, this should probably be replaced with the solid metal blocks that were used to hold the QWPs. If these blocks are reinstalled, the waveplate mount should be twisted slightly in yaw in order to reduce the amount of backscatter into the cavities.

  1536   Tue Dec 16 21:31:21 2014 KateDailyProgress Minor changes to layout of cavity transmitted path

Evan, Kate

We made a few minor modifications to the optical breadboard in transmission of the cavities.

For one, we removed the QWPs which were the first optics in the transmission paths. These had been necessary for the prior cavities where the Silica Tantala mirror coatings were not birefringent. The circular polarization which was transmitted needed to be turned into linear polarization to get the beat note on the PD. Now, because the cavities with AlGaAs coatings are birefringent, the resonant and transmitted light is already linearly polarized and the QWPs unnecessary. Before removing them, the power on the main readout PD, a PD1811, was 208 mV. Afterwards, it was 194 mV. 

Second, we started to set up the fiber coupler to send some light to the ATF lab where it will later be used in a PLL to stabilize the laser used for the seismometer sensing. There's a 29.5 uW pick-off of the North cavity transmitted light which had been dumped. I found a high reflector mirror to put in its place to direct light to the fiber. I also made sure the fiber coupler at the other end is secured to the table and the output dumped. A first attempt to couple the light did not work, but I need to find a way either to monito remotely the power transmitted or just temporarily feed the fiber back into the CTN lab. 

 

  1535   Tue Dec 16 14:06:35 2014 Aidan, Kate, EvanSummaryEnvironmentTask list

Some tasks not included on the list:

  • Temperature loops
    • Cavities
    • Can
    • PMC
  • Noise candidates:
    • Scatter
    • RFAM
    • PDH loop noise
Attachment 1: algaasTasks.jpg
algaasTasks.jpg
  1534   Mon Dec 15 20:43:49 2014 Aidan, Kate, EvanDailyProgressPMCSouth Faraday isolator needs help

Quote:

The laser beam entering the first Faraday isolator appears to be 1–2 mm too low. It is clipping on the input aperture, and the transmitted beam looks like crap.

When Aidan and I turned on the south laser today, we found that the transmitted beam out of this Faraday was entirely crap. It was blindingly obvious on an IR card, and only 50 uW was making it to the input of the PMC. The rest was scattering at wide angles at the Faraday output port.

It is not clear to me how the pointing through the Faraday could have deteriorated, since it is on a solid metal mount and is only 10 cm from the output of the laser.

At any rate, I was able to "recover" the previous performance (i.e., crappy but workable) by placing the Faraday isolator slightly further down in the optical path. Before, the layout was:

Laser -> QWP -> HWP -> Faraday -> lens -> HWP -> steering mirror -> PMC EOM,

and the HWP angles were -1 deg and 167 deg, respectively.  Now the layout is

Laser -> QWP -> HWP -> lens -> steering mirror -> Faraday -> HWP -> PMC EOM,

and the HWP angles are 341 deg and 167 deg. The first HWP angle is chosen so that 20 mW is transmitted through the Faraday (the rest is dumped at the Faraday's various output ports). The second HWP angle is chosen to send s polarization through the PMC EOM. I then had to resteer through the PMC EOM and through the PMC. With 20 mW incident on the PMC, the transmission is 11 mW. Not great, but about the same as the previous situation.

I remark that the south optical path between the laser and the PMC should be reworked as soon as is feasible, because what I've done is a hack job to keep things moving. Either the Faraday mount needs to be remachined, or the optical path needs to be redesigned to allow for proper steering through the Faraday. Additionally, the table surface next to the laser mounts is noticeably warm to the touch, so I do not recommend trying to shim up the laser (as it may negatively impact the heatsinking).

  1533   Wed Dec 10 23:59:54 2014 EvanNotesNoiseBudgetThe sense in which phi_c is a "coating loss angle"

I thought I had posted this several months ago, but I cannot find it now.

I believe this document explains how the "coating loss angle" ϕ_c (as measured in an optical experiment) is related to the true mechanical loss angles of the coating materials (as measured by a ringdown). In general they aren't the same, if I'm understanding the formalisms of Nakgawa and Hong correctly.

Attachment 1: hongbreakout.pdf
hongbreakout.pdf hongbreakout.pdf
  1532   Sun Nov 16 23:32:49 2014 EvanDailyProgressopticNorth photothermal TF

I believe the factor of π / F here is an error. It should instead be the transmission T. That lowers the absorption estimates to more like 5 ppm.

 

  1531   Sun Nov 9 18:54:15 2014 EvanSummaryDocumentationHow to run the CTN experiment

A manual for running the CTN experiment is attached. I'll update and expand as needed.

Attachment 1: Manual.pdf
Manual.pdf
  1530   Fri Oct 3 12:32:00 2014 EvanDailyProgressFSSSouth TTFSS input-referred noise

Yesterday I took some TFs and noise spectra on the south TTFSS with the loop open and the beam blocked. Relevant information:

  • Gains were 800 common, 800 fast
  • Excitations were injected into COM EXC
  • The relevant test points I monitored were COM TP 4 (henceforth "com"), FAST OUT 2 (henceforth "fast"), and HV TP 4 (henceforth "HV").
  • At each test point I took a TF with the excitation on, and a noise spectrum with the excitation off.
  • I also took a TF from COM EXC to COM OUT 2, so that I can use the known gain of the first amplifier (−4 V/V) to refer everything to COM OUT 1.

Then using the noise spectra, I divided by the relevant TF to arrive at an input noise referred to COM OUT 1.

The results are attached. The low/high frequency upswings on the HV trace are due to the input noise of the SR785.

Since the common, fast, and HV traces all lie on top of each other, I interpret this to mean that the noise of all of them is dominated by sources occurring upstream of COM TP 4. So the TTFSS is limited by the noise of its input stages, with a spectrum (at COM OUT 1) of 25(5) nV/rtHz.

I also measured the slope of the south PDH error signal (with 3 mW incident, and with fresh mode-matching), and found a slope of 8.4 V/MHz, as measured at COM OUT 1. This gives a frequency noise of 3.0(6) mHz/rtHz, which is well below the current beat level.

Attachment 1: inputNoise.pdf
inputNoise.pdf
Attachment 2: southInputNoise.zip
  1529   Tue Sep 30 21:45:14 2014 EvanDailyProgressPMCRe-inserted PMC, moved BB EOM

Quote:

I will optimize the alignment through these components, check the polarization, and then take a new beat.

I found that the beat above 100 Hz was about 10 times worse than before, with a 2 kHz hump similar to what I saw on the north cavity before I separated its BB and resonant EOMs.

I suspect this is some kind of effect involving light bouncing back and forth several times between the two EOMs.

To remedy this, I took out the post-PMC Faraday isolator and put the BB EOM in its place. This gives a longer path length between the EOMs. Then I realigned through the EOMs and took a beat. Now I've recovered the 0.03–0.05 Hz/rtHz level that I had yesterday morning.

I turned up the incident powers on the cavities to 3 mW, and then optimized the mode-matching. However, I do not seem to be able to push down the beat any further. So perhaps it is now limited by something else.

I tried reinserting the Faraday isolator between the two EOMs, but could not place it in such a way to get the beam to transmit through. Since it's not an essential component, I think I'm going to leave it out for the time being rather than undertake a huge realignment marathon.

  1528   Mon Sep 29 23:28:37 2014 EvanDailyProgressPMCRe-inserted PMC, moved BB EOM

I swept the south laser with a triangle wave and optimized the mode-matching as best I could using the periscope mirrors and the translation stages. I got to a visibility of 0.3, which stinks (the maximum is 0.7 with these birefringent coatings).

I took a beat spectrum (attached) and noticed that the noise around 0.1–1 kHz is improved. Indeed, by reducing the visibility south I find the beat gets worse.

I decided some more involved mode-matching (involving beam profiling and alm simulation) is needed.

Before setting up the beam profiler, I noticed that the beam entering the cavity does not appear Gaussian, as seen on an IR card.

The laser beam entering the first Faraday isolator appears to be 1–2 mm too low. It is clipping on the input aperture, and the transmitted beam looks like crap.

Neither the Faraday nor the laser itself have any alignment adjustment knobs. I therefore had to choose between two evils: shim up the laser mount (and thereby risk having to realign the entire optical path, as well as possibly reducing the heatsinking of the mount to the table) or reinsert the PMC and move the BB EOM.

I opted for the latter: I reinserted the PMC, removed the EOAM (+QWP+PBS), placed the resonant EOM where the EOAM used to be, and then placed the BB EOM where the resonant EOM used to be.

I will optimize the alignment through these components, check the polarization, and then take a new beat.

Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1527   Thu Sep 25 16:50:09 2014 EvanDailyProgressFSSShould probably move south BB EOM

We've noticed for a while now that we cannot turn up the gain on the south TTFSS as high as on the north TTFSS, despite having similar optical power levels, similar mode-matching, etc. (See the OLTFs in ctn:1504.) The north gain can be set to 900/900 on the common/fast knobpots, but on south it's more like 600/600.

Because the BB EOM is placed before the PMC, I suspect the cavity pole of the PMC (1.8 MHz, measured in elog: in 2010) is giving us extra phase which prevents us from turning the loop gain up higher. Indeed, when I remove the PMC from the south optical path (and realign into the south cavity) I find I can turn the south TTFSS knobpots up to 800/800. A new OLTF is probably in order.

The easy thing to do for now is to leave the PMC out. The better thing is probably to move the BB EOM to come after the PMC. Since there's no room, this probably means putting the BB EOM where the resonant EOM currently is, put the resonant EOM where the EOAM currently is, and then put the EOAM elsewhere. The EOAM could just as well come before the PMC, since we're only attempting intensity stabilization well below the PMC cavity pole.

  1526   Tue Sep 23 18:40:08 2014 EvanDailyProgressEOMNo more 2 kHz hump in beat

Background

Yesterday I think I narrowed down the source of the 2 kHz frequency noise hump: it is voltage noise from the TTFSS being injected into the broadband EOM.

With the north cavity unlocked (and the TTFSS set to "test"), I monitored the (undemodulated) RAM using the auxiliary 1811 and the HP4395A. There were clear, broad 600 Hz humps on either side of the 14.75 MHz carrier. It disappeared when I unplugged the drive to the broadband EOM.

Then I looked at various test points on the TTFSS HV board with the SR785. On the COM → EOM path, the TF shaping takes the COM noise and produces (what I think is) the same 600 Hz bump, which is then sent to the EOM. In the beat, the bump appears at 2 kHz because of the north TTFSS boost; with the boost off, it reverts to 600 Hz.

This is the case on both TTFSS boards, but it only leaked into the beat on the north cavity. So I suspected it was an issue with how the EOMs are aligned on the north path. On north, the BB EOM was immediately followed by the resonant PDH EOM; on south, between the BB EOM and PDH EOM there is a PMC, an FI, and some other optics.

Today's work

I moved the resonant EOM so that it follows the EOAM. After the post-EOAM PBS, I did the following:

  • I set down a HWP, and then used a temporary PBS to ensure s-polarization of the beam.
  • A few inches after the first HWP, I set down a second HWP and used a temporary PBS to ensure p-polarization of the beam.
  • Between the HWPs, I placed the resonant EOM, screwed it down, and then aligned the beam through it.

Then I redid the mode-matching into the north cavity and measured the beat. I kept it locked for about 90 minutes and didn't see the 2 kHz hump appear, so I'm guessing this solved the issue.

To do

  • Minimize RAM on north cavity
  • RIN data is stale and needs to be retaken
  • Need to fix a nominal operating power for beat PD (I pick 7 dBm, because we're using a ZRPD-1 phase detector)
  • Marconi noise data is stale and needs to be retaken
  • PLL readout data is stale and needs to be retaken
  • Seismic data is stale and needs to be retaken
Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1525   Fri Sep 19 12:01:13 2014 EvanDailyProgressEnvironmentNorth EOM heater

I added a 48 Ω kapton heater to the north resonant EOM. It's got 40 mA going through it right now; no loop yet.

  1524   Thu Sep 18 22:13:35 2014 EvanDailyProgressRFAMRAM coherence

I used the auxiliary 1811 as an out-of-loop RAM monitor. The RF from the 1811 is mixed with the PDH LO, and then low-passed at 1.9 MHz.

I'm not sure about the RAM calibration here. I took the raw spectrum (in V/rtHz), multiplied by 10^(4/20) (assuming 4 dB conversion loss in the mixer), then divided by the measured dc voltage (about 20 mV), then divided by 40 (because of the different dc/ac tranimpedances).

Anyway, the point is that the 200 Hz hump we see in the beat seems to be from the north RAM.

Attachment 1: ramCoherence.zip
Attachment 2: ramCoherence.pdf
ramCoherence.pdf
  1523   Thu Sep 18 13:57:00 2014 EvanNotesISSEOAM phenomenology

I've been unsure of how the EOAMs are affecting the state of the light impinging on the cavity.

So far we've been rotating the post-EOAM QWPs so as to maximize the strength of the amplitude modulation. I'm still not sure what this does. I'd like to instead fix the QWP at ±45° and then insert the EOAM, regardless of whether this introduces a DC power offset. At the very least this will give us a polarization state that we think we understand.

Attachment 1: eoam.pdf
eoam.pdf eoam.pdf
  1522   Wed Sep 17 17:47:39 2014 EvanDailyProgressFSSPDH block diagram

Since the straightforward tabletop optimizations (mode-matching, RAM minimization) have not been able to make the high-frequency excess beat noise disappear, perhaps it is time to undertake a more systematic study of the PDH loop noise and add these traces to the noise budget.

Here's my interpretation of the PDH block diagram for one of the two cavities.

Attachment 1: ctnBlock.pdf
ctnBlock.pdf
  1521   Tue Sep 16 15:08:57 2014 EvanDailyProgressBEATAttempts at scatter reduction

Quote:

I went through the table today looking for ghost beams. Most were already dumped. For those that weren't, I put down a dump or an iris.

I again looked at TTFSS OUT2 with the cavities unlocked (i.e., the open-loop error signals) and found that the low-frequency seismic/scatter wall appears only on south. So I hunted around south for a while. I found a series of ghost beams reflecting off the EOAM input and hitting dangerously close to the EOM output aperture. So I moved the EOAM forward a few inches, then adjusted its kinematic mount to offset these beams a bit. The EOAM should be realigned, and we should check to make sure the ghost beams are not entering the EOM again.

With the increased space between the EOM and EOAM, I installed a flipper mirror that takes the beam to the 1811. Then I minimized the RAM (from –54 dBm to –72 dBm with 85 mV dc).

FM dev: 10 kHz

Averages: 10, 50, 100, 500

I added a flipper mirror before the north EOAM, as well as a HWP before the resonant EOM (so that we can independently control the polarization entering the two EOMs). I optimized the RAM, but saw no improvement in the beat.

  1520   Mon Sep 15 18:36:32 2014 taraNotesVacuumtransmission of vacuum windows

 We measured the transmission of the vacuum windows. The total transmission through two windows is 0.975 +/- 0.002.

 

  •   This measurement is for checking how much the power is transmitted through the vacuum windows. It can be used for calibration of photo thermal measurement for better accuracy of absorption estimation.
  •  The beam is incident normal to the two windows. The power before the vac tank and after the vac tank were measured. The incident level was set ~ 4 mW. I used a Thorlab power meter with ND filter off to measured the power.
  •  
  Pin [mW] Pout [mW]
1 3.9 +/- 0.02 3.81+/- 0.02
2 3.96 3.87
3 3.98 3.88
4 3.99 3.88
  Tavg 0.975

 

If we assume that both windows have the same transmission, the transmission for each window will be 0.988.

  1519   Mon Sep 15 18:29:47 2014 EvanDailyProgressBEATAttempts at scatter reduction

I went through the table today looking for ghost beams. Most were already dumped. For those that weren't, I put down a dump or an iris.

I again looked at TTFSS OUT2 with the cavities unlocked (i.e., the open-loop error signals) and found that the low-frequency seismic/scatter wall appears only on south. So I hunted around south for a while. I found a series of ghost beams reflecting off the EOAM input and hitting dangerously close to the EOM output aperture. So I moved the EOAM forward a few inches, then adjusted its kinematic mount to offset these beams a bit. The EOAM should be realigned, and we should check to make sure the ghost beams are not entering the EOM again.

With the increased space between the EOM and EOAM, I installed a flipper mirror that takes the beam to the 1811. Then I minimized the RAM (from –54 dBm to –72 dBm with 85 mV dc).

FM dev: 10 kHz

Averages: 10, 50, 100, 500

  1518   Mon Sep 15 07:18:38 2014 EvanNotesBEATnote for tonight beat

Quote:

 RCAV transPD_DC :0.54 V

ACAV transPD_DC: 0.16 V (loop might oscillate when DC level was measured, need to double check)

 In this noise budget I corrected the PDH shot noise level (incident power is 1 mW, visibility is 0.5).

Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1517   Mon Sep 15 04:32:21 2014 taraNotesBEATnote for tonight beat

 RCAV transPD_DC :0.54 V

ACAV transPD_DC: 0.16 V (loop might oscillate when DC level was measured, need to double check)

  1516   Mon Sep 15 02:24:48 2014 taraNotesTempCtrlepic for slow control is disabled

Perl scripts for controlling the vacuum tank and slow feedback to the two lasers are acting weird. Usually we can run three scripts simultaneously, but I just notice that only two can be run at the same time. When I restarted the script for acav feedback, the vac temp control stopped. When I restarted the vac temp control, acav slow feedback stopped.  I'll check this later.

  1515   Sun Sep 14 14:46:44 2014 EvanDailyProgressBEATReinstalled aux 1811

I reinstalled the auxiliary 1811 on the input side of the table.

I tried to get the free-running noise of the north cavity by locking south and then using the PLL to read out the beat from the 1811. However, even on a FM deviation of 400 kHz / Vrms, I could not get the PLL to lock. So I suspect the free-running noise is just too high to use this method.

I also used this aux 1811 to optimize the PDH EOM alignments. I saw no change in the beat spectrum after doing this. I would like to demodulate the 1811 signal using the PDH LO, but this will require some reconfiguration of the RF distribution.

  1514   Sun Sep 14 10:37:57 2014 EvanDailyProgressBEATNew beat

Quote:

We turned on both ISS loops today.

Here is an in-loop characterization of the south RIN with and without the ISS.

Here is the beat measurement from Thursday, with the ISS loops on and the table floated.

I've fudged the photothermal noise slightly by just using twice the south cavity's PT measurement, rather than south and north. I need to take RIN data from north, and then I can add south + north PT.

Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1513   Thu Sep 11 21:43:32 2014 Tara, EvanDailyProgressISSSouth ISS

We turned on both ISS loops today.

Here is an in-loop characterization of the south RIN with and without the ISS.

Attachment 1: RINsouth.pdf
RINsouth.pdf
Attachment 2: southRIN.zip
  1512   Thu Sep 11 11:40:41 2014 EvanDailyProgressopticSouth photothermal TF

Quote:

I took a swept-sine measurement of the photothermal TF just as Tara and I did for the north cavity. To get a better measurement, I made some configuration changes:

  • I turned the power incident on the south cavity up to 8.5 mW by adjusting the post-laser HWP from 318° to 286°.
  • I placed an OD2.0 in front of the beat PD to prevent RF saturation.

Settings/values:

  • The beat was at 13.8 MHz.
  • The PLL Marconi was on 50 kHz FM deviation, and the SR560 gain was 100 V/V.
  • South transmission PD was 460(5) mV dc.
  • South transmission power (directly out of vacuum chamber) was 2.20(5) mW dc.

The results are attached. I'm not sure why there's a discrepancy around 200 Hz between the two traces. Below 100 Hz the measurement looks relatively clean.

The light rejected out of the post-EOAM PBS is only 2 mW (compared with 9 mW transmitted), which makes me suspicious that the post-EOAM QWP is not rotated properly, or else the input polarization into the EOAM is wrong. We should check this before redoing this measurement.

As with the north cavity, I find that an absorption of 6 ppm is needed make the measured curve lie on top of the theory curve.

For the time being, I have left the input power at 8 mW in case we want to take this again tomorrow. There's currently a dump upstream of the PMC to block the beam.

I didn't like the EOAM situation, so I rotated the post-EOAM QWP from 302° to 285°. With no voltage applied to the EOAM, this gives 6 mW of p and 6 mW of s. This may not be the true optimal setting, but the previous 2 mW / 9 mW situation seems too weird to be right. For commissioning the ISS I suspect we'll have to redo this EOAM setup to make sure the polarizations are behaving as we think.

The results are attached. I'm still seeing discrepancies at the points where the TFs are stitched together. Maybe it's because I'm using the SR785's auto source adjust feature.

 

Attachment 1: spt.pdf
spt.pdf
Attachment 2: southPT.zip
Attachment 3: photothermalTF.pdf
photothermalTF.pdf
  1511   Thu Sep 11 00:47:14 2014 EvanDailyProgressopticSouth photothermal TF

I took a swept-sine measurement of the photothermal TF just as Tara and I did for the north cavity. To get a better measurement, I made some configuration changes:

  • I turned the power incident on the south cavity up to 8.5 mW by adjusting the post-laser HWP from 318° to 286°.
  • I placed an OD2.0 in front of the beat PD to prevent RF saturation.

Settings/values:

  • The beat was at 13.8 MHz.
  • The PLL Marconi was on 50 kHz FM deviation, and the SR560 gain was 100 V/V.
  • South transmission PD was 460(5) mV dc.
  • South transmission power (directly out of vacuum chamber) was 2.20(5) mW dc.

The results are attached. I'm not sure why there's a discrepancy around 200 Hz between the two traces. Below 100 Hz the measurement looks relatively clean.

The light rejected out of the post-EOAM PBS is only 2 mW (compared with 9 mW transmitted), which makes me suspicious that the post-EOAM QWP is not rotated properly, or else the input polarization into the EOAM is wrong. We should check this before redoing this measurement.

As with the north cavity, I find that an absorption of 6 ppm is needed make the measured curve lie on top of the theory curve.

For the time being, I have left the input power at 8 mW in case we want to take this again tomorrow. There's currently a dump upstream of the PMC to block the beam.

Attachment 1: spt.pdf
spt.pdf
Attachment 2: southPT.zip
Attachment 3: photothermalTF.pdf
photothermalTF.pdf
  1510   Tue Sep 9 23:22:41 2014 Tara, EvanDailyProgressFSSSouth 200 kHz oscillation; cavity pole

We locked the south cavity using the north TTFSS. 200 kHz oscillation is still present, so whatever this is probably doesn't reside in the TTFSS box.

Tara and I also took cavity pole measurements using the EOAM and two PDA10CSs, one placed before the cavity (monitoring the light rejected from the post-EOAM PBS) and one after the cavity (on the ISS breadboard). The HP4395A was used to drive the EOAM, and we then took the transfer function which takes the pre-cavity PD voltage to the transmission PD voltage. I will fit these later, but the poles appear to be consistent with the values tabulated in ctn:1475.

Attachment 1: algaasFinesse.pdf
algaasFinesse.pdf
Attachment 2: algaasFinesse.zip
  1509   Tue Sep 9 12:22:31 2014 EvanDailyProgressBEATReduced south scatter shelf

Last night, I looked at the TTFSS OUT2 on the spectrum analyzer, with the cavity unlocked and the laser PZT and broadband EOM unplugged. I believe this should give the demodulated spectrum of the RAM.

For north, the spectrum was white. For south, the spectrum was white down to 10 Hz, and below 10 Hz there was an apparent scattering shelf that rose several orders of magnitude above the white spectrum.

Tara and I went through the south path and placed an OD0.3 ND filter between various optics. We tried twisting one of the mode-matching lenses to reduce the scatter shelf, but it didn't seem to work.

This morning I went through the south path again with an OD0.5 filter, and eventually focused on the lens right after the FI. I found that I could greatly reduce the height of the scattering shelf (relative to the height of the white noise) by placing the filter downstream of lens, but not upstream of it. So I twisted this lens slightly, re-modematched into the cavity (minimal adjustment was required), and I found that the scatter shelf was reduced to less than 1 order of magnitude above the white noise.

Then I took a beat spectrum (carrier at 11.5 MHz, seems stable). The scatter shelf in the beat is not reduced by much, so I'll have to think about where to look next.

Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1508   Mon Sep 8 12:41:32 2014 EvanDailyProgressTempCtrlChamber temp change

The beat is currently at 450 kHz. So I've changed the setpoint on the chamber temperature from 31.200 °C to 31.300 °C. We'll see if this pushes the beat to a higher frequency.

Edit: actually I decided to leave the chamber temperature alone and instead adjust the north cavity heater again. With a DVM, I measured the initial heater voltage as 10.64 V. I turned the power supply knob ever so slightly to get 10.60 V instead.

  1507   Mon Sep 8 04:20:41 2014 ranaDailyProgressFSSTTFSS OLTFs, residual frequency noise

 

The AD829 cannot drive 100 Ohms rail to rail, so you can only go into the 50 Ohm input devices if the signals are small (less than 1-2 V I guess).

  1506   Sun Sep 7 20:37:50 2014 Tara, EvanDailyProgressopticSome lens tweaks

Some minor maintenance/improvements to the optical setup:

  • We replaced the existing lens before the beat PD (RoC = 51.5 mm) with a slightly faster lens (RoC = 38 mm) in order to reduce the spot size on the diode.
  • Tara improved the clamping of mode-matching lenses before the south cavity (they weren't tightened down enough before)
  1505   Sun Sep 7 19:33:40 2014 Tara, EvanDailyProgressopticNorth photothermal TF

Tara and I took another photothermal TF of the north cavity today. Relevant parameters:

  • Power incident on cavity: 10 mW (up from the usual 1 mW)
  • Beat frequency: 1.2 MHz, drifting to 650 kHz (we are hoping it will swing through 0 Hz overnight and settle above a few megahertz by tomorrow)
  • DC voltage on north ISS PD: 2.39(5) V
  • DC power transmitted through cavity: 3.77(2) mW
  • PLL actuation coefficient: 50 kHzpk / 1 Vrms
  • PLL UGF: 80 kHz (measured)
  • EOAM drive: 5 Vpp from 20 kHz to 300 Hz, then 3 Vpp from 300 Hz to 0.2 Hz

In the attached data I have already converted the raw data (in V/V) into hertz of beat frequency per watt of circulating power. For this I use the conversion factor (50 kHz / 21/2 V) × (2.39 V / 3.77 mW) × π / F, with F = 16 700. Since the TF (again) appears to be junk above 1 kHz, I haven't bothered undoing the CLTF of the PLL.

The attached plot shows the expected photothermal TF in terms of hertz of beat frequency per watt of absorbed power per mirror. Therefore, the scaling factor that makes our measurement (given in hertz per watt of circulating power) overlap with the expected TF (given in hertz per watt of absorbed power per mirror) should be the average absorption of each mirror. I find that this scaling factor is 6 ppm, which seems surprisingly low, especially given our earlier finding that we have at least 120 ppm of scatter + absorption loss. So I will double check for missing factors of 2, 4, π, etc.

At any rate, the shape of the measured transfer function appears to be in good agreement with the expectation up to 100 Hz. If we believe that the coating/substrate photothermal crossover happens around 10 Hz, and we believe our measurement from 10 Hz to 100 Hz, then this seems to indicate that the thermo-optic cancellation has been somewhat successful.

Attachment 1: photothermalTF.pdf
photothermalTF.pdf
Attachment 2: northPT.zip
  1504   Sun Sep 7 11:54:21 2014 EvanDailyProgressFSSTTFSS OLTFs, residual frequency noise

Acutally it does look like it's a 50 Ω loading issue. I find that when 50 Ω inline terminators are added to OUT1 and OUT2, the measured OLTF is reduced by a factor of 1.6. This explains the discrepancy between the SR785 and HP4395 measurements. I've attached the corrected OLTF plots, along with plots of a vector fit, and the expected residual frequency noise [assuming a free-running NPRO noise of 104 Hz/Hz1/2 × (1 Hz / f)].

South UGF is at 200 kHz with almost no phase margin. We need to fix this.

Attachment 1: pdhTFs.pdf
pdhTFs.pdf
Attachment 2: oltfNorth.pdf
oltfNorth.pdf
Attachment 3: oltfSouth.pdf
oltfSouth.pdf
Attachment 4: frnoise.pdf
frnoise.pdf
Attachment 5: noiseBudget.pdf
noiseBudget.pdf
  1503   Sat Sep 6 12:54:05 2014 Tara, EvanDailyProgressopticNorth photothermal TF

Tara and I took an SR785 measurement of the north photothermal transfer function.

Clearly there's something wrong with the measurement above 1 kHz.

Attachment 1: nPT.pdf
nPT.pdf
Attachment 2: npttf.zip
  1502   Thu Sep 4 19:39:36 2014 Tara, EvanDailyProgressFSSTTFSS OLTFs

With the cavities locked, Tara and I took OLTFs of the PDH loops.

Below 100 kHz, we used the SR785 with a 70.7 mVpk excitation. Above 30 kHz, we used the HP4395A with a 22.4 mVrms excitation (these are the "HF" traces on the attached plot).

Before taking these TFs, we turned the loop gains up as high as possible without making the loops saturate.

  • For north, the gains were 900 common and 900 fast, and the incident power was 1.26(5) mW.
  • For south, the gains were 632 common and 770 fast, and the incident power was 1.05(2) mW.

We injected the excitation on common EXC. We then measured the TF which takes common OUT2 to common OUT1. This is almost the OLTF of the loop, except for an AD829 between OUT2 and OUT1 which has a gain of −4 V/V.

Currently I'm not sure how to explain the magnitude discrepancy between the SR785 and HP4395 measurements. Both OUT1 and OUT2 have a 50 Ω output impedance, so I would expect the impedance difference between the SR785 and HP4395 would cancel out in this measurement.

Attachment 1: pdhTFs.pdf
pdhTFs.pdf
ELOG V3.1.3-